Figure 1 compares the sensitivity with the beam shifted vertically by 6mm on WE (blue), at the mirror center (red) and before the WE replacement (yellow).
- Displacing the beam spot by 6mm allows to retrieve the same optical gain as before the WE replacement and consequently the same shot noise. https://logbook.virgo-gw.eu/virgo/?r=66626
- There is a comb of lines spaced by approximately 3Hz that is present between 10Hz and 100Hz. This is not due to a point absorber. It could be due to the WE vacuum pump that is still in operation, but it can be something else that got damaged during the WE replacement.
- There is a new bump at 178Hz
- There is a broadband noise between 50Hz and 300Hz. It is this noise that makes us loose 15 Mpc. It may be related to the comb or to the bump. It level changes when the losses are reduced by moving the beam spot on WE, but less so than shot noise, most of the apparent change is in any case due to the reduction in shot noise level when the losses are reduced.
In summary there are 4 issues in the sensitivity curve. Point 1) is clearly related to a point absorber like issue on the WE mirror, the rest of the points are most likely due to something else, point 4) is the one that costs us the most Mpc.